# IterPref: Focal Preference Learning for Code Generation via Iterative Debugging

### v2

@@ 5 Main Results
Table 1 presents a comparison between baseline models, DPO variants, and IterPref. We discuss the findings from the following perspectives.
IterPref Achieves Fine-grained Alignment.
We begin by noting that the preference pairs generated through the iterative process exhibit unique characteristics distinct from those found in other preference sets, such as Code-Optimise. These characteristics are reflected in the performance decline observed when directly applying vanilla DPO or RPO algorithms. While certain configurations, such as DS-Coder-7B-Instruct-DPO and Qwen2.5-Coder-7B-Base RPO, demonstrate performance gains on the full BigCodeBench set, it remains a consistent trend that DPO and RPO generally underperform relative to the baseline models. Typically, when an external LLM attempts to correct an error in a code, only a small portion of the given code is modified to fix the logic or runtime error, while most of the code remains unchanged. This results in highly similar preference pairs. Overly similar preference pairs, where identical tokens appear in both positive and negative examples, introduce ambiguity and noise to the policy model, thereby weakening alignment gains.
This performance degradation highlights the need for explicit measures to guide the policy model in focusing on critical tokens that drive functional errors, preventing it from being misled by identical parts. IterPref addresses this by masking identical tokens in the dispreferred code and explicitly contrasting the critical tokens extracted, which results in significant performance gains. As shown in Table 1, IterPref-DPO outperforms DPO by an impressive 3.3% and 5.9% on average across benchmarks, with IterPref-RPO achieving gains ranging from 6.3% to 12.4%. Compared to Code-Optimise, IterPref achieves significant performance gains over both instruction-tuned Code LLMs and base models, with a 4.0% improvement using DS-Coder-7B-Instruct and a 7.4% improvement using Qwen2.5-Coder-7B, respectively.
IterPref Achieves Significant Gains over Existing Alignment Frameworks.
The preference pair construction used by Code-Optimise involves sampling multiple solutions and testing for a given instruction. It is a straightforward and effective method that demonstrates strong performance, as shown in Figure 5. However, constructing pairs based solely on pass rates fails to capture how erroreous code should be corrected, thus limiting the model’s ability to generalize improvements. Compared to baselines AceCoder and Code-Optimise, IterPref consistently achieves performance gains over both, using the same base Code LLM.
Figure 5: Average results on HumanEval, MBPP, and BigCodeBench-Full based on Qwen2.5-Coder-7B.
IterPref Improves Challenging Coding Tasks.
@@
IterPref drives performance gains not only in basic coding tasks but also in more challenging ones. We highlight that the IterPref framework has the potential to boost Code LLMs to solve complex coding tasks. Notably, Qwen2.5-Coder-7B equipped with IterPref achieves a 29.7% pass@1 score on BigCodeBench Complete Hard, matching the performance of larger Code LLMs DeepSeek-Coder-V2-Instruct (29.7) and Claude-3-Opus (29.7) Anthropic (2024), and approaching Llama-3.1-405B-Instruct (30.4) Grattafiori et al. (2024). When given more attempts, IterPref achieves pass@5 of 45.7%, outperforming DeepSeek-R1 (40.5)  DeepSeek-AI et al. (2025) and GPT-o1 (40.2). On the Instruct Hard split, pass@5 of the IterPref-Qwen is 34.7%, comparable to the performance of GPT-o3-mini (33.1).
6 Analysis
In this section, we present a statistical analysis of common failure case types to pinpoint frequent pitfalls in code generation. Additionally, we compare the costs of generating and annotating preference pairs to guide more efficient preference alignment and reduce errors.
IterPref Generates Fewer Errors.
By contrasting critical tokens between a corrected version and its preceding iteration explicitly, a Code LLM equipped with IterPref makes fewer errors. Figure 6 illustrates the frequency of common failure types in the BigCodeBench Complete-Full set, showing a notable reduction in error occurrences when using IterPref-RPO compared to RPO and Code-Optimise on Qwen2.5-Coder-7B. This suggests that while RPO includes SFT, it still requires targeted learning of critical errors in the dispreferred samples to effectively reduce mistakes.
IterPref Provides an Efficient Pathway for Preference Annotation.
We compare the cost of synthesizing one preference pair between IterPref and the sampling techniques adopted by Code-Optimise, primarily considering external LLM calls and execution times. Given an instruction, Code-Optimise synthesizes $m$ code snippet candidates (where $m$ is often set to 100), using $n$ test cases from the raw dataset, leading to $m\times n$ executions on the CPU. In contrast, for a single instruction, IterPref requires up to 7 LLM calls and executions. Considering the failure ratio (when code can’t pass the generated test cases), an estimated 10.4 calls are needed for a given instruction, which is far fewer than the practice of the current method. Though massive sampling is diverse and effective, it is not efficient as most code snippets are discarded. IterPref shows that starting with a single code snippet, even if it fails initially, it still holds the potential to form a preference pair for alignment training.
Figure 6: The frequency of the most common failure types on the BigCodeBench Complete-Full set.
@@ 7 Ablation Study
Despite the effectiveness of IterPref in pinpointing critical error regions, there remain open questions about how best to incorporate negative examples and how much context is truly beneficial for code correction. We therefore explore several settings: (i) SFT: Supervised fine-tuning using the positive sample from the preference pair; (ii) Hybrid Training: Half of the samples in a batch are trained using vanilla DPO, while the other half follows the IterPref approach; (iii) Diff-Augmentation: provide more context for the dispreferred sample by including 1 or 2 lines of tokens before and after $D^{-}$; and (iv) Symmetric Masking Strategy: The Code LLMs learn from the tokens in $D^{+}$ rather than the full sequence of positive sample. We conduct experiments on CodeQwen1.5-7B-Chat and base model Qwen2.5-Coder-7B, with the results presented in Table 2.
Hybrid Training & Diff-Augmentation both expose Code LLMs to dispreferred samples but differ in scope: Hybrid Training uses the entire dispreferred sequence, while Diff-Augmentation focuses on a small token window around the changed segment. Although adding extra tokens around $D^{-}$ may seem beneficial, it often confuses the model and lowers performance. Concentrating on only the most critical tokens leads to better outcomes, highlighting the importance of accurately identifying them for IterPref. The iterative debugging process naturally suits this approach because only a limited portion of the code typically changes between iterations, leaving the bulk of the code intact. Consequently, it produces samples in which the differences can be isolated with high precision. Although extracting $D^{-}$ with the Longest Common Subsequence (LCS) is not always perfect, it reliably captures the key modifications and thus narrows the search space for crucial tokens. Crucially, this automated method obviates the need for manual annotations, underscoring its efficiency and practicality.
Supervised Fine-Tuning (SFT) ranks as the second-best approach, following the IterPref framework. Because IterPref employs iterative debugging, it ensures that the preferred sample in each preference pair is validated to pass test cases, thereby guaranteeing its high quality. Consequently, straightforward implementation of SFT also yields notable performance gains. However, the SFT process disregards the dispreferred sample, missing the chance to learn from common mistakes that should be avoided. In contrast, IterPref not only increases the model’s likelihood of producing correct tokens by fully leveraging error-free code, but also systematically highlights and penalizes the specific tokens that lead to critical errors. By explicitly learning from accurate coding patterns while avoiding common mistakes, IterPref achieves more fine-grained alignment and ultimately outperforms standard SFT.
Setting HumanEval MBPP Avg.

### v3

@@ 4 Main Results
Table 1 presents a comparison between baseline models, DPO variants, and Target-DPO. We discuss the findings from the following perspectives.
Focal Alignment Outperforms Global Alignment. Preference pairs from iterative debugging differ significantly from those in datasets like Code-Optimise, causing a performance drop with vanilla DPO or RPO. While some settings, like DS-Coder-7B-Instruct-DPO, show gains on BigCodeBench, DPO and RPO generally underperform compared to baselines. In typical correction scenarios, an LLM modifies only a small portion of the code to fix errors, creating highly similar preference pairs. This overlap introduces ambiguity, as identical tokens in both positive and negative examples weaken the model’s ability to distinguish meaningful differences.
This degradation underscores the need for explicit mechanisms to focus the policy model on tokens responsible for functional faults. Target-DPO addresses this by emphasizing error tokens in the dispreferred code and explicitly contrasting the critical edits. As shown in Table1, Target-DPO outperforms DPO by 3.3% and 5.9% on average across benchmarks, while Target-RPO yields improvements ranging from 6.3% to 12.4%.
Figure 4: Comparison with CodeDPO, PLUM, and Code-Optimise using DeepSeekCoder-6.7B. Additional results are provided in Appendix B.3.
Target-DPO Achieves Significant Improvements over Methods that Rely on Coarse-grained Pass Rate Signals. While methods like PLUM and CodeDPO, which construct preference pairs by testing multiple sampled solutions, offer a straightforward and effective approach, their reliance on coarse-grained pass/fail signals inherently limits the model’s ability to learn nuanced error correction and generalize improvements. As shown in Figure 4, this limitation becomes apparent when compared to our Target-DPO.
@@
Target-DPO Improves Challenging Coding Task. We highlight that the Target-DPO framework has the potential to boost Code LLMs to solve complex coding tasks. Notably, Qwen2.5-Coder-7B equipped with Target-DPO achieves a 29.7% pass@1 score on BigCodeBench Complete Hard, matching the performance of larger Code LLMs DeepSeek-Coder-V2-Instruct (29.7%) and Claude-3-Opus (29.7%) Anthropic (2024), and approaching Llama-3.1-405B-Instruct (30.4%) Grattafiori et al. (2024). When given more attempts, Target-DPO achieves pass@5 of 45.7%, outperforming DeepSeek-R1 (40.5%)  DeepSeek-AI et al. (2025) and GPT-o1 (40.2%). On the Instruct Hard split, pass@5 of the Target-DPO-Qwen is 34.7%, comparable to the performance of GPT-o3-mini (33.1%).
Figure 5: Illustration for Target-DPO and its ablations. Target-DPO rewards correct code tokens while penalizing only error-specific tokens in rejected code, teaching models to truly understand code through targeted alignment.
@@ 5 Ablation Study
HumanEval(Avg) MBPP(Avg) BCB(Complete) BCB(Instruct) Average
SFT (EpiCoder 40k) 83.9 75.5 50.9 39.1 62.4
SFT (EpiCoder 80k) 85.1 78.9 52.3 39.4 63.9
SFT (EpiCoder 380k) 85.7 77.8 53.4 43.8 65.2
SFT (CodeFlow 59k) 85.5 75.7 51.6 39.1 62.9
Our Target-DPO (59k) 87.2 76.5 53.3 43.1 65.0
Table 2: Results of EpiCoder-SFT with varying amounts of training data and our method on Qwen2.5-Coder-7B.
Aug Hybrid HumanEval MBPP BCB-Inst BCB-Comp Average
Base Plus Base Plus
CodeQwen1.5-7B-Chat - - 83.5 78.7 79.4 69.0 39.6 43.6 65.6
Target-DPO 83.5 79.3 81.5 66.1 38.2 44.5 65.5
83.5 76.8 80.2 65.1 34.7 44.5 64.1
89.6 85.4 83.9 69.8 39.9 48.7 69.6
Target-RPO 84.8 79.3 81.5 67.7 34.6 44.0 65.3
86.0 79.9 81.5 65.9 36.0 45.0 65.7
89.6 86.0 82.5 70.4 38.3 48.4 69.2
Table 3: Ablation study on how much contextual information from negative examples is beneficial for Target-DPO, evaluated using CodeQwen1.5-7B-Chat. Additional results with Qwen2.5-Coder-7B are provided in Appendix C.
HumanEval(Avg) MBPP(Avg) BCB(Complete) BCB(Instruct) Avg.
SFT (Correct) 85.5 75.7 51.6 39.1 63.0
SFT (Incorrect) 82.7 73.5 49.2 38.6 61.0
Table 4: Performance comparison of SFT on correct and incorrect code from CodeFlow uisng Qwen2.5-Coder-7B.
@@
Despite the effectiveness of Target-DPO in pinpointing critical error regions, there remain open questions about how best to incorporate negative examples and how much context is truly beneficial for code correction. We therefore explore several settings: (i) SFT: Supervised fine-tuning using the positive sample from the preference pair; (ii) Hybrid Training: Half of the samples in a batch are trained using vanilla DPO, while the other half follows the Target-DPO approach; (iii) Diff-Augmentation: provide more context for the dispreferred sample by including 1 or 2 lines of tokens before and after $D^{-}$; and (iv) Symmetric Masking Strategy: The Code LLMs learn from the tokens in $D^{+}$ rather than the full sequence of positive sample. In Figure 5, we illustrate these settings.
Supervised Fine-Tuning. A comparison with EpiCoder-SFT, considering varying amounts of training data, is shown in Table 2. Our Target-DPO achieves performance comparable to the strong SFT baseline EpiCoder-380k using only 59k training samples (about one-sixth), unveiling the power of targeted alignment.
The positive samples undergo iterative debugging and are validated by test cases, they maintain high quality, allowing SFT to achieve reasonably strong performance. However, SFT overlooks dispreferred samples, missing the opportunity to contrast and precisely align positive and negative examples. In contrast, Target-DPO not only leverages error-free code to increase the likelihood of correct code but also precisely penalizes tokens responsible for critical errors, achieving finer-grained alignment and better performance consequently.
Hybrid Training & Diff-Augmentation. Both settings expose Code LLMs to more tokens from the dispreferred samples but differ in scope: in Hybrid Training, 50% of the training samples use the entire dispreferred sequence, while Diff-Augmentation provides a small token window around the $D^{-}$. Table 3 shows that while adding extra context around $D^{-}$ may appear beneficial, it often introduces noise that confuses the model, making it unclear which parts need local alignment, ultimately leading to degraded performance.
We find that concentrating solely on the most critical tokens yields better results, highlighting the importance of accurately grounding these tokens for more effective targeted alignment. The iterative debugging process naturally supports this precise localization, as typically only a small portion of the code changes between iterations, while the majority remains unchanged. These targeted regions can be easily identified using the Longest Common Subsequence (LCS), allowing meaningful differences to be isolated with high precision.
Symmetric Masking Strategy. When training with the symmetric masking, where Code LLMs learn from both $D^{+}$ and $D^{-}$ without access to the full positive sample, the model struggles to retain its core code generation capabilities and fails to benchmark effectively. The primary goal of Code LLMs is to generate complete and correct code. Although learning symmetrically from both $D^{+}$ and $D^{-}$ may seem appealing, the focus should be on ensuring Code LLMs learn from fully correct code rather than fragmented pieces. Without complete code contexts, the positive sample cannot properly align with the instruction, leading to incomplete and misleading signals in the learning process.
Generated Test Cases can Distinguish Good and Error Code. Table 4 compares supervised fine-tuning using either preferred or dispreferred samples. The results show that SFT on preferred samples outperforms that on dispreferred ones by an average of 2.0%. This quality gap between pairs, introduced through debugging iterations, suggests that test cases effectively differentiate high- and low-quality code snippets, providing training pairs with clear quality contrast. In our debugging pattern, the quality differences between code snippets are primarily influenced by the feedback from test cases, indicating that test cases can reliably distinguish between good and bad code when verified through dedicated efforts.
6 Related Work
Code Language Models. Powerful Code LLMs like Qwen2.5-Coder Hui et al. (2024a), DeepSeek-Coder Guo et al. (2024), StarCoder Li et al. (2023); Lozhkov et al. (2024), Magicoder Wei et al. (2024) and EpiCoder Wang et al. (2025b) demonstrate their capabilities in various code generation tasks. Current Code LLMs primarily focus on supervised fine-tuning during the post-training stage. While SFT enables Code LLMs to learn the correct patterns, it fails to effectively make them aware of incorrect patterns or how to rectify errors in code. In this work, Target-DPO framework aims to enable Code LLMs to further learn through pairwise contrasting of critical tokens Lin et al. (2024), allowing Code LLMs to continually improve.
Preference Pair Construction. Existing methods construct preference pairs by ranking candidate solutions based on pass rates. PLUM Zhang et al. (2024) constructs preference pairs by ranking candidate code solutions based on passed test cases. Code-Optimise Gee et al. (2025) incorporates efficiency as an additional learning signal, augmented with annotations from unit test feedback and execution time. AceCoder Li et al. (2024) selects pairs with distinct pass rate differences. DSTC Liu et al. (2024b) constructs preference pairs using self-generated code and tests. A related concurrent work is CodeDPO, which formulates preference learning as a direct optimization problem using pass/fail signals and proposes a PageRank-inspired algorithm to select high-quality preference pairs. In contrast, our method aligns code LLMs through error-resolving edits rather than relying on coarse-grained execution outcomes. This fine-grained supervision provides richer training signals that better capture the semantics of code correction, resulting in improved performance, as verified in Figure 4.
7 Conclusion
We present Target-DPO, a novel preference alignment framework that emulates human iterative debugging to capture critical errors in incorrect code for precise optimization. Target-DPO first identifies error-prone regions and applies an improved DPO algorithm contrasting pivotal segments, teaching Code LLMs to understand and correct code through targeted preference alignment, achieving promising coding performance. To support this framework, we efficiently synthesize high-quality preference pairs to create CodeFlow, where each sample undergoes iterative refinement until it passes unit tests, with the modification history providing a natural record of error corrections. Extensive experiments show that Target-DPO-equipped Code LLMs achieve significant performance improvements in code generation and excel in tackling basic and complex coding tasks.
Limitations
Target-DPO is inspired by the debugging pattern of developers, serving as a novel framework for fine-grained preference learning in Code LLMs. Instead of using pass rate alone, Target-DPO derives preference pairs from iterative debugging process. By contrasting critical tokens between a corrected version and its preceding iteration, Target-DPO helps the model to understand code through targeted alignment. However, this study focuses on a dataset of 59k samples without further expansion, which may limit generalizability, but offers opportunities for future exploration with larger data.
Acknowledgments
This work was partly supported by the National Natural Science Foundation of China (Grant No. 62576191), the research grant No. CT20240905126002 of the Doubao Large Model Fund, and the National Natural Science Foundation of China (No. 62306216).
Appendix
In this appendix, we first provide more details of our core methodology, including preference pair construction and implementation specifics (Section A). Section B then introduces the evaluation benchmarks and presents comprehensive experimental results, showcasing performance on various benchmarks and detailed comparisons against EpiCoder and other relevant methods. Subsequently, we provide in-depth analyses such as scaling laws, ablation studies on key parameters, data diversity assessments, error pattern examinations, and efficiency evaluations (Section C).
Appendix A Methodology and Data Construction
This section details the core methodology of our proposed approach, including the iterative refinement process for preference pair construction and the generation and quality assessment of synthetic test data. Implementation specifics relevant to these methodological aspects are also covered.
A.1 Iterative Refinement and Preference Pair Construction
The core of our data generation relies on an iterative refinement process. Figure 6 illustrates the progression of code sample pass rates through successive refinement iterations using execution verification feedback.
As depicted in Figure 6, at the first attempt (iter0), only 36.7% of the code samples pass their corresponding test file, indicating that debugging is necessary for the remaining code. Failed codes go through continual refinement, with the pass rate gradually approaching 67.5%. The pass rate rises sharply from iter 1 to iter 3 and then slows. Between iter 4 and iter 5, only 1.2% of cases improve, indicating incremental benefits from further iterations. Thus, additional iterations are not considered.
Current methods construct preference pairs based on pass rate signals and conduct DPO to optimize Code LLMs. Two notable limitations arise: one from the data and one from the algorithm.
Regarding preference data, a snippet with a low pass rate may only require minor modifications to become correct, as errors tend to be isolated to specific parts of the code. We address this by iterative debugging with editing traces naturally annotated by the differences between iterations.
Regarding the algorithm, relying solely on full preference learning can introduce noise during optimization, as positive and negative pairs can be highly similar. This not only hinders the model from learning more effective error correction patterns but also increases the risk of overfitting. We solve this by explicitly identifying which parts of the code need to be aligned.
Iterative debugging can pass through 67.5% of tasks within a 5-time API call budget for each task. But given just 5 sampling attempts, the pass rate falls to 51.90% averaged across 5k samples, as shown in Table 5. This initial comparison highlights that iterative debugging can achieve a higher pass rate under similar API constraints.
API Calls Pass rate (%)
Debugging (Ours) Up to 5 times 67.5
Sampling 5 times for each task 51.9
Table 5: Pass Rate of Debugging and Sampling for Preference Pair Construction.
To further investigate the limits of sampling for difficult cases, we collected 5k samples that didn’t succeed within 5 debugging iterations. Table 6 details the pass rate when applying N sampling solutions to these difficult tasks.
Sampled Solutions N Passed Failed
5 2.42% 97.58%
10 3.26% 96.26%
15 3.96% 96.04%
30 4.20% 95.80%
50 4.36% 95.64%
Table 6: Pass Rate of Sampling for Difficult Cases (Failed within 5 Debugging Iterations).
The results in Table 6 indicate that for samples which could not succeed with 5 iterations using interpreter feedback and runtime error information, additional sampling alone yields very low pass rates (e.g., only 3.96% pass with 15 sampling attempts). This suggests such cases can hardly pass through additional sampling alone.
To directly compare the effectiveness of preference pairs generated via iterative debugging versus sampling, we conducted experiments using 10k training samples. Table 7 presents these comparative results.
HumanEval HumanEval+ MBPP MBPP+ BCB-Inst BCB-Comp LCB-v5 Avg.
Ref. 61.6 53.0 76.9 62.9 40.2 45.8 24.1 52.1
Sampling 84.1 79.3 82.3 68.5 40.8 49.2 32.4 62.3
Debugging 87.8 84.8 83.3 69.7 41.1 49.6 32.5 64.1
Table 7: Results of Preference Pair Construction using Sampling vs. Debugging (10k training samples).
As shown in Table 7, iterative debugging can generate more meaningful preference pairs than sampling by leveraging interpreter feedback and runtime information, thereby achieving better results (e.g., an average score of 64.1 vs 62.3) with lower API costs.
A.2 Synthetic Test Case Generation
A.2.1 Rationale for using Synthetic Test Cases
We address the relationale behind systhetic test cases from the following perpectives. Optimizing code LLMs through preference learning requires a large amount of training data, which is difficult to annotate or verify manually. Synthetic data has become a widely adopted approach. For example, Qwen2.5-Coder utilizes tens of millions of synthetic instruction samples, and models like DeepSeek-V3 and R1 also incorporate synthetic data during training, also as demonstrated in studies like PLUM, SelfCodeAlign and DSTC.
A.2.2 Validity of Synthetic Test Cases
We have made the following efforts to ensure the qaulity of test cases: (i) First, we adopted a powerful LLM, GPT-4o, as the test case generator to primarily ensure its validity. (ii) Through prompting engineering, we have invested significant effort into making the generated test cases broad and meaningful. (iii) We conducted a manual evaluation by performing a random sample check. We manually examined 100 data samples and found that all the generated test cases correctly reflected the task requirements. However, we observed that these test cases tend to be relatively simple and may not cover all edge cases.
A.2.3 Coverage Analysis of Test Cases
To validate the effectiveness of test cases in exercising source code, we conducted coverage analysis on a sample of 1,000 training instances. Code coverage, a crucial metric in software testing, quantifies the extent to which a program’s source code is exercised by test cases. This metric measures the percentage of code executed by a test suite, which we evaluated using the Python Coverage library [^1] https://github.com/nedbat/coveragepy. Our evaluation framework organizes the source code and test cases into directories, maintaining a clear separation between the source code (implemented as Python modules) and the corresponding unit tests (developed using the unittest [^2] https://docs.python.org/3/library/unittest.html framework). The coverage metric is calculated through the following equation:
[…]
The results of the coverage analysis are shown in Figure 7.
Figure 7: Coverage Distribution of Test Case.
A.2.4 Evaluating the quality of test cases using LLM-as-a-judge
To evaluate test case quality, we employ the LLM-as-a-judge approach, assessing three dimensions: accuracy, effectiveness, and reasonableness on a 5-point scale. Detailed evaluation prompts are provided in 9. We sample 1,000 data points from the training data and conduct evaluations using the DeepSeek-V3-0324 model. The evaluation results in Figure 8 demonstrate satisfactory test case quality across all dimensions.
Figure 8: Comparative distribution of LLM-as-a-judge scores across three evaluation metrics: accuracy, effectiveness, and reasonableness.
Figure 9: Prompt for evaluating the quality of test cases using LLM-as-a-judge.
A.2.5 Prompt and Examples for Generated Test Cases
To illustrate our synthetic test case generation process, Figure 10 displays the prompt template provided to the LLM. Following this prompt, Figure 11 presents an example of the test cases generated for question “Create a Python function named ‘generate_arithmetic_sequence’ that generates the first N terms of an arithmetic sequence given the first term and the common difference. The function should take three parameters: the first term (a), the common difference (d), and the number of terms (N). The function should return a list containing the first N terms of the sequence”.
Figure 10: Prompt used for generating code and test cases.
Figure 11: An example of generated test cases.
A.3 Training and Inference Parameters
Unless specified otherwise for a particular experiment, our code LLMs were trained with consistent hyperparameter settings. The learning rate was set to $1\times 10^{-5}$ for the 7B code LLMs and $5\times 10^{-6}$ for the 14B models. We used a global batch size of 128, full-parameter training 3 epochs with max prompt length of 1024 and generation length of 2048. A cosine learning rate scheduler with was employed, with 3% of the total training steps dedicated to warm-up. For the DPO algorithm, $\beta$ is set to 0.1, and $\alpha$ is set to 1.0 for RPO. $\pi_{\theta}$ and $\pi_{\text{ref}}$ are both initialized with the weights of the evaluated model, while $\pi_{\text{ref}}$ keeps frozen during training. For all inference, greedy decoding was utilized with pass rate at first attempt reported. For Code-Optimise, we replicated their setup, and the $DPO_{PvF}$ setting results are reported. For PLUM and CodeDPO, their scores are taken from their respective papers.
Appendix B Experimental Results and Comparisons
This section begins by describing the evaluation benchmarks and their statistics. It then presents the comprehensive experimental results of our method on these benchmarks, followed by detailed comparisons against Supervised Fine-Tuning (SFT) and other relevant state-of-the-art methods.
B.1 Evaluation Benchmarks: Description and Statistics
We detail the individual function-level code generation benchmarks used for evaluation in this subsection. Table 8 summarizes key statistics for these benchmarks, such as the number of problems and the average number of tests per problem.
Dataset Problems Avg. Tests
HumanEval 164 9.57
HumanEval+ 748.07
MBPP 378 3.11
MBPP+ 105.40
Easy   279 18.07
LiveCodeBench Medium 331 21.81
Hard   270 24.78
Table 8: Statistics of Evaluation Benchmarks.
HumanEval and MBPP are popular benchmarks for assessing code generation. Considering the limited test cases in these benchmarks (HumanEval: 9.57 avg. tests; MBPP: 3.11 avg. tests, as seen in Table 8), we followed previous work and utilized the EvalPlus framework to evaluate model robustness across a broader range of test cases (HumanEval+: 748.07 avg. tests; MBPP+: 105.40 avg. tests). To ensure fair comparison, we used version 0.2.0 of MBPP+ provided by EvalPlus [^3] https://github.com/evalplus/evalplus v0.3.1 , which removes some broken tasks (399 $\to$ 378 tasks).
BigCodeBench (BCB) is a comprehensive benchmark designed to assess a model’s ability to handle real-world programming tasks, particularly its effectiveness in utilizing various function calls as tools. Our model’s ability to adeptly manage these high-complexity scenarios underscores its suitability for BigCodeBench.
LiveCodeBench (LCB), statistics for which are also included in Table 8 (showing problem distribution by difficulty), is a benchmark designed to evaluate code generation models on challenging competitive programming problems, often sourced from real coding contests. Unlike benchmarks focused solely on function completion, LCB tasks typically require more complex algorithmic reasoning and problem-solving skills. The evaluation often simulates a contest environment, potentially including hidden test cases to assess the robustness and correctness of the generated solutions under pressure. In our experiments, we utilize LiveCodeBench-v5 (LCB-v5) to gauge the model’s capabilities in tackling these demanding, contest-style coding scenarios.
B.2 Comparison with Supervised Fine-Tuning
We compare our method with standard Supervised Fine-Tuning (SFT) to demonstrate the benefits of our preference alignment approach. The overall comparison of our method (Target-RPO) against SFT and reference models on various benchmarks is presented in Table 9. For instance, on Qwen2.5-Coder-7B, our Target-RPO achieves an average of 65.3, surpassing the SFT baseline’s 63.2.
HumanEval MBPP BCB-Inst BCB-Comp LCB-v5 Average